OpenStudy (anonymous):

A fair coin is flipped three times. What is the probability of getting heads at least twice? Write your answer as a simplified fraction.

OpenStudy (anonymous):

You can use a Bernoulli Trial for this problem. \[\left(\begin{matrix}n \\ r\end{matrix}\right)=p ^{r}\times q ^{n-r}\] n is the amount of trials (coin flips) = 3 r is the result (2 OR 3) = 2/3 p is the probability of getting a head = 1/2 r is the probability of getting a tail = 1/2 We will have to do this twice, once for 2 heads and once for 3. So we get: \[\left(\begin{matrix}3 \\ 2\end{matrix}\right)\times (1/2)^{2} \times (1/2)^{3-2}\] 3 x 1/4 x 1/2 = 3/8 \[\left(\begin{matrix}3 \\ 3\end{matrix}\right)\times (1/2)^{3} \times (1/2)^{3-3}\] 1 x 1/8 x 1 = 1/8 3/8 + 1/8 = 1/2 So the probability of getting heads at least twice is 1/2

OpenStudy (anonymous):

@tfitz17 perfect solution :)